find such that y=x^2-16x+c has exactly one real root
Answer by stanbon(75887)

find such that y=x^2-16x+c has exactly one real root
------
a = 1 ; b = -16 ; c
-------
Solve: b^2 - 4*1*c = 0
----
256 - 4c = 0
4c = 256
c = 64
======================
Cheers,
Stan H.